# Density-Aware Farthest Point Sampling

### v1

@@ 8 Numerical Experiments
We focus on molecular property prediction, a regression task in quantum chemistry. We use the QM7, QM8, and QM9 datasets, which contain 7,165, 21,766, and 130,202 molecules, respectively. Our setup is based on Climaco & Garcke (2024), but with important differences. While Climaco & Garcke (2024) studied FPS mainly for small training sets (1% to 10% of the data), we evaluate DA-FPS for larger training set sizes, from 5% up to 20% of the data pool. This allows us to address challenges with high-dimensional density estimation in very small data regimes (which DA-FPS requires), and to test DA-FPS in scenarios where FPS alone does not significantly outperform standard passive sampling in terms of average prediction error. Our data preprocessing is similar to Climaco & Garcke (2024), but we reduce the feature vector dimension (up to 276 instead of up to 1300) to further mitigate issues with high-dimensional density estimation. Full details on datasets, preprocessing, and descriptors are in Appendix G. Appendix H provides empirical evidence that the datasets satisfy the required data assumption B.2. We use two regression models that have been utilized in previous works for molecular property prediction tasks: kernel ridge regression with a Gaussian kernel (KRR) (Stuke et al., 2019; Deringer et al., 2021) and feed-forward neural networks (FNNs) (Pinheiro et al., 2020). Both models are Lipschitz continuous, which is required for our theoretical results (see Appendix I for details on the models and hyperparameter tuning). We evaluate prediction performance using the mean absolute error (MAE), defined as $\text{MAE}:=\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}|y_{i}-\tilde{y}_{i}|$, where $y_{i}$ are true values, $\tilde{y}_{i}$ are predictions, and $n_{u}$ is the number of unlabeled points in the data pool. Let us remark that the final goal of our experiments is to empirically show the benefits of using DA-FPS compared to other model-agnostic state-of-the-art sampling approaches and investigate the benefits of complementing classical passive sampling approaches with the FPS. We do not make any claims on the general prediction quality of the employed models on any of the studied datasets.
@@
Baseline sampling strategies
We compare our approach with four sampling techniques, among those analyzed in well-established papers related to data selection (Sener & Savarese, 2018; Ghorbani & Zou, 2019; Mirzasoleiman et al., 2020; Killamsetty et al., 2021), that fit our application scenario and problem constraints: passive sampling model-agnostic data selection techniques that do not rely on the knowledge of the labels. Specifically, we consider random sampling (RDM), Farthest Point Sampling (FPS), facility location (Frieze, 1974) and $k$-medoids++ (Mannor et al., 2011). Both, facility location and $k$-medoids++, attempt to minimize a sum of pairwise squared distances. However, the fundamental difference is that facility location is a greedy technique, while $k$-medoids++ is based on a segmentation of the data points into clusters.
@@
It is important to note that, as explained in Section 7, during the initial phase of the sampling process DA-FPS uses uniform weights. Meaning that, DA-FPS initially coincide with FPS. The amounts of points selected with uniform weights is controlled by the hyperparameter $u$. In Appendix L we empirically investigate the benefits of combining FPS with the baselines used to benchmark DA-FPS. The experiments in Appendix L show that augmenting the baselines approaches with FPS during the initial steps of the sampling process consistently leads to a decrease in the MAE of the predictions. Still, DA-FPS remains the most competitive approach. This investigation highlights that the advantages of DA-FPS stem not only from initially using FPS but also from its density-aware weighting, taking place after the first $u$ points have been selected.
Regarding the DA-FPS hyperparameter $u$, determining the amount of points selected with uniform weights, we make observations that can guide a user to a suitable choice. Our ablation study on the ZINC dataset (Appendix M) shows that for larger training set sizes ($10\%$), the choice of $u$ has little effect on DA-FPS performance compared to the baselines. However, for smaller training set sizes, $u$ has a stronger impact. When $u$ is large, DA-FPS behaves more like FPS, which is known to reduce the maximum absolute prediction error (MAXAE $\max\limits_{1\leq i\leq n_{u}}|y_{i}-\tilde{y}_{i}|$), but may not lead to better average error. The work in Climaco & Garcke (2024) found that, on the QM datasets, selecting just the first 2% of points with FPS, then switching to random sampling, achieves similar MAXAE as using FPS alone. This suggests that $u$ should not be chosen too large (e.g., $5\%$ of the available data points). Based on these observations we think that an “optimal” choice $u$ depends on what the user wants to achieve. Qualitatively, if the user is interested in ensuring low MAXAE, at the cost of higher MAE, a larger $u$ is recommended. If the user is mainly interested in performing well on average in the region of high data density, a smaller $u$ is recommended. In Appendix M, we also analyze how changing the DA-FPS hyperparameter $k$ affects performance. $k$ defines the amount of $k$-nearest neighbors considered for computing the weights. The results show that DA-FPS works well for a range of $k$ values, but choosing $k$ that is too small can reduce its effectiveness.
@@ 9 Conclusion
We proposed DA-FPS, a passive model-agnostic sampling algorithm to select training sets by weighted fill distance minimization. We provided an upper bound for the approximation error of DA-FPS to the estimated weighted fill distance minimization problem we defined. Our numerical results demonstrated that utilizing DA-FPS to select training sets has a positive impact on the average prediction of the regression models, particularly for larger datasets, in line with our theoretical motivation. Concerning the choice of DA-FPS hyperparameter $u$, we hypothesize that statistical properties, such as density or size of the tails of the dataset, may be exploited by a user to take an informed decision. Future research directions should investigate alternatives for the data density estimations. In this paper, we used a $k$-nearest neighbors (kNN) approach for density estimation. Using the kNN was computationally efficient but lead to scaling issues when estimating the weight function used in the definition of the weighed fill distance. Alternative methods that overcome this issue while remaining effective and computationally efficient could be developed.
@@
Selecting points aiming to reduce the training sets weighted fill distance is beneficial when traditional labeling methods, such as numerical simulations or laboratory experiments, are expensive or time-consuming. In such applications ML regression models provide fast label predictions for new data points. The quality of these predictions depends on the quality of the training data. Therefore, selecting an effective training set is crucial to ensure accurate predictions. Our research on minimizing the training set weighted fill distance can potentially identify sets that can enhance average prediction quality for various regression models and tasks, preventing the wastage of expensive labeling resources on training sets that may only benefit a specific learning model or task.
@@ Appendix A The weighted fill distance is non-negative
The weighed fill distance defined in (2) is non-negative.
@@ Appendix B Assumptions
We recapture two assumptions from Climaco & Garcke (2024) that underlie our theoretical results. The first assumption concerns the data being analyzed and the relationship between data features and labels.
@@
With (17) we assume that the expected error on the training set is bounded. Moreover, with the Lipschitz continuity assumptions we limit our study to error functions that show a certain regularity. However, these regularity assumptions on the error function are not too restrictive and are connected with the regularity of the evaluated trained model as already shown in Climaco & Garcke (2024). For instance, the $\lambda_{l_{\mathcal{Y}}}$-Lipschitz regularity and the convexity in the second argument are verified by all $L_{p}$-norm error functions, with $1\leq p<\infty$. We also assume that $\mathbb{E}[|l({\bm{x}},Y,m_{\mathcal{L}})|\big{|}{\bm{x}}]<\infty$ for any given ${\bm{x}}\in\mathcal{X}$. This assumption formalizes the intuitive fact that, in applications, independently of the trained model and the feature vector considered, we can expect the error function to take finite values.
@@
Note that in (18) the expectation $\mathbb{E}\left[l({\bm{x}},Y,m_{\mathcal{L}})|{\bm{x}}\right]$ is independent of $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$. This is because, as mentioned in Section 3, we consider a scenario where $p_{\mathcal{D}}$ may differ from $p_{\mathcal{L}}$ but the map connecting a data location ${\bm{x}}\in\mathcal{X}$ and its associated label value is independent on how the data is selected, that is, $p_{\mathcal{D}}\neq p_{\mathcal{L}}$ and $p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. In what follows we define $p(y|{\bm{x}}):=p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. Next, fixed ${\bm{X}}=\tilde{{\bm{x}}}\in\mathcal{X}$, we want to bound $\mathbb{E}\left[l(\tilde{{\bm{x}}},Y,m_{\mathcal{L}})|\tilde{{\bm{x}}}\right$]. To do that we use a result from Climaco & Garcke (2024): For fixed $\tilde{{\bm{x}}}\in\mathcal{X}$ and ${\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}$ we have
@@
$+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})\int_{\mathcal{X}}p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$\leq$ $\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}>0}{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}}{\int}\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}C\left(1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}\right)p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$\leq$ $\;C\thinspace\sup_{{\bm{x}}\in\mathcal{X}}\left(\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}\left(1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}\right)\right)+\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}}+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$=$ $\;CW_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)+\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}}+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right].$
The last inequality follows from taking the supremum over ${\bm{x}}\in\mathcal{X}$, taking the constant terms out of the integral and noticing that $\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}>0}{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}}{\int}p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}\leq\underset{\mathcal{X}}{\int}p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}=1$. ∎

### v2

@@ 8 Numerical Experiments
In this section we empirically validate DA-FPS on molecular property prediction tasks. We describe our experimental setup, introduce the baseline sampling strategies and regression models, and analyze performance using mean absolute error as our primary evaluation metric.
@@
We focus on molecular property prediction, a regression task in quantum chemistry. We use the QM7, QM8, and QM9 datasets, which contain 7,165, 21,766, and 130,202 molecules, respectively. Our setup is based on Climaco and Garcke (2024), but with important differences. While Climaco and Garcke (2024) studied FPS mainly for small training sets (1% to 10% of the data), we evaluate DA-FPS for larger training set sizes, from 5% up to 20% of the data pool. This allows us to address challenges with high-dimensional density estimation in very small data regimes (which DA-FPS requires), and to test DA-FPS in scenarios where FPS alone does not significantly outperform standard passive sampling in terms of average prediction error. Our data preprocessing is similar to Climaco and Garcke (2024), but we reduce the feature vector dimension (up to 276 instead of up to 1300) to further mitigate issues with high-dimensional density estimation. Full details on datasets, preprocessing, and descriptors are in Appendix G. Appendix H provides empirical evidence that the datasets satisfy the required data assumption B.2. We use two regression models that have been utilized in previous works for molecular property prediction tasks: kernel ridge regression with a Gaussian kernel (KRR) (Stuke et al., 2019; Deringer et al., 2021) and feed-forward neural networks (FNNs) (Pinheiro et al., 2020). Both models are Lipschitz continuous, which is required for our theoretical results (see Appendix I for details on the models and hyperparameter tuning). We evaluate prediction performance using the mean absolute error (MAE), defined as $\text{MAE}:=\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}|y_{i}-\tilde{y}_{i}|$, where $y_{i}$ are true values, $\tilde{y}_{i}$ are predictions, and $n_{u}$ is the number of unlabeled points in the data pool. The code necessary to reproduce the results presented in this section is available in our GitHub repository. [^2] https://github.com/PaClimaco/DA-FPS Let us remark that the final goal of our experiments is to empirically show the benefits of using DA-FPS compared to other model-agnostic state-of-the-art sampling approaches and investigate the benefits of complementing classical passive sampling approaches with the FPS. We do not make any claims on the general prediction quality of the employed models on any of the studied datasets.
Baseline sampling strategies
We compare our approach with four sampling techniques, among those analyzed in well-established papers related to data selection (Sener and Savarese, 2018; Ghorbani and Zou, 2019; Mirzasoleiman et al., 2020; Killamsetty et al., 2021), that fit our application scenario and problem constraints: passive sampling model-agnostic data selection techniques that do not rely on the knowledge of the labels. Specifically, we consider random sampling (RDM), Farthest Point Sampling (FPS), facility location (Frieze, 1974) and $k$-medoids++ (Mannor et al., 2011). Both facility location and $k$-medoids++ attempt to minimize a sum of pairwise squared distances. However, the fundamental difference is that facility location is a greedy technique, while $k$-medoids++ is based on a segmentation of the data points into clusters.
@@
It is important to note that, as explained in Section 7, during the initial phase of the sampling process DA-FPS uses uniform weights. Meaning that, DA-FPS initially coincide with FPS. The amount of points selected with uniform weights is controlled by the hyperparameter $u$. In Appendix L we empirically investigate the benefits of combining FPS with the baselines used to benchmark DA-FPS. The experiments in Appendix L show that augmenting the baselines approaches with FPS during the initial steps of the sampling process consistently leads to a decrease in the MAE of the predictions. Still, DA-FPS remains the most competitive approach. This investigation highlights that the advantages of DA-FPS stem not only from initially using FPS but also from its density-aware weighting, taking place after the first $u$ points have been selected.
Regarding the DA-FPS hyperparameter $u$, determining the amount of points selected with uniform weights, we make observations that can guide a user to a suitable choice. Our ablation study on the ZINC dataset (Appendix M) shows that for larger training set sizes ($10\%$), the choice of $u$ has little effect on DA-FPS performance compared to the baselines. However, for smaller training set sizes, $u$ has a stronger impact. When $u$ is large, DA-FPS behaves more like FPS, which is known to reduce the maximum absolute prediction error (MAXAE $\max\limits_{1\leq i\leq n_{u}}|y_{i}-\tilde{y}_{i}|$), but may not lead to better average error. The work in Climaco and Garcke (2024) found that, on the QM datasets, selecting just the first 2% of points with FPS, then switching to random sampling, achieves similar MAXAE as using FPS alone. This suggests that $u$ should not be chosen too large (e.g., $5\%$ of the available data points). Based on these observations we think that an “optimal” choice $u$ depends on what the user wants to achieve. Qualitatively, if the user is interested in ensuring low MAXAE, at the cost of higher MAE, a larger $u$ is recommended. If the user is mainly interested in performing well on average in the region of high data density, a smaller $u$ is recommended.
In Appendix M, we also analyze how changing the DA-FPS hyperparameter $k$ affects performance. $k$ defines the amount of $k$-nearest neighbors considered for computing the weights. The results show that DA-FPS works well for a range of $k$ values, but choosing $k$ that is too small can reduce its effectiveness. The rough magnitude of $k$ likely depends on the data dimension and distribution and is generally investigated in the context of density estimation. The problem of choosing $k$ is common for kkNN-based density estimation methods. We expect further insights from that domain. We think that the choice of $k$ should take into account at least two aspects: First, $k$ should not be too small, because it would lead to high variance in the values of the weights, and consequently, to noticeable differences in the values of weights associated with points lying in regions with similar true density. Second, $k$ should not be too large as the computational cost of DAFPS depends linearly on $k$.
@@ 9 Conclusion
We proposed DA-FPS, a passive model-agnostic sampling algorithm to select training sets by weighted fill distance minimization. We provided an upper bound for the approximation error of DA-FPS to the estimated weighted fill distance minimization problem we defined. Our numerical results demonstrated that utilizing DA-FPS to select training sets has a positive impact on the average prediction of the regression models, particularly for larger datasets, in line with our theoretical motivation.
A key aspect of DA-FPS is the need to estimate the data densities. Poor density estimation in DA-FPS can lead to ineffective weights that cause oversampling of high-density regions, generating redundant information in the selected set, and undersampling of low-density regions, leading to underrepresentation of important areas of the data space in the training set. This can lead to degraded model performance, for instance, through increased prediction errors in underrepresented regions. In this paper, we used a $k$-nearest neighbors (kNN) approach for density estimation. Using the kNN was computationally efficient but led to scaling issues when estimating the weight function used in the definition of the weighted fill distance. Alternative methods that overcome this issue while remaining effective and computationally efficient could be developed. Our ablation study on the hyperparameter $k$ in Appendix M provides empirical evidence that choosing $k$ too small can negatively impact performance, highlighting the importance of appropriate density estimation. Future research directions should investigate alternatives for the data density estimations or approaches for direct density ratio estimation. Concerning the choice of DA-FPS hyperparameter $u$, we hypothesize that statistical properties, such as density or size of the tails of the dataset, may be exploited by a user to take an informed decision. Developing adaptive strategies for automatically selecting the hyperparameters $k$ and $u$ based on dataset characteristics would enhance the robustness and applicability of DA-FPS.
Notation
Symbol Description
$\mathcal{X}$ Feature space, subset of $\mathbb{R}^{d}$
$\mathcal{Y}$ Label space, subset of $\mathbb{R}$
$\mathcal{D}$ Pool of available data points, $\{({\bm{x}}_{i},y_{i})\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$
$\mathcal{D}_{\mathcal{X}}$ Set of feature vectors of points in $\mathcal{D}$, $\{{\bm{x}}_{i}\}_{i=1}^{n}\subset\mathcal{X}$
$\mathcal{L}$ Training set $\{({\bm{x}}_{i_{j}},y_{i_{j}})\}_{j=1}^{b}\subset\mathcal{D}$, with $i_{j}\in\{1,\ldots,n\}$
$\mathcal{L}_{\mathcal{X}}$ Set of feature vectors of points in $\mathcal{L}$, $\{{\bm{x}}_{i_{j}}\}_{j=1}^{b}\subset\mathcal{D}_{\mathcal{X}}$
$p_{\mathcal{D}}$ Data source distribution on $\mathcal{X}\times\mathcal{Y}$
$p_{\mathcal{L}}$ Training data distribution on $\mathcal{X}\times\mathcal{Y}$
$p_{\mathcal{X}_{\mathcal{D}}}$ Marginal of $p_{\mathcal{D}}$ on $\mathcal{X}$
$p_{\mathcal{X}_{\mathcal{L}}}$ Marginal of $p_{\mathcal{L}}$ on $\mathcal{X}$
$\hat{p}_{\mathcal{X}_{\mathcal{D}}}$ General density estimator of $p_{\mathcal{X}_{\mathcal{D}}}$
$\hat{p}_{\mathcal{X}_{\mathcal{L}}}$ General density estimator of $p_{\mathcal{X}_{\mathcal{L}}}$
$\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}}}$ $k$-NN density estimator of $p_{\mathcal{X}_{\mathcal{D}}}$
$\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}}}$ $k$-NN density estimator of $p_{\mathcal{X}_{\mathcal{L}}}$
$\psi_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ Weight function: $1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}$
$W_{\mathcal{L}_{\mathcal{X}},\mathcal{X}}(p_{\mathcal{X}_{\mathcal{L}}}\|p_{\mathcal{X}_{\mathcal{D}}})$ Weighted fill distance of $\mathcal{L}_{\mathcal{X}}$ in $\mathcal{X}$
$W^{k}_{\mathcal{L}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}}}$ Estimated weighted fill distance using $k$-NN
$r^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ Adaptive neighborhood radius for density estimation
$\omega^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$ Number of points in $k$-neighborhood of $\bm{x}$} within a distance $r^{k}_{\mathcal{L}_{\mathcal{X}}}({\bm{x}})$
$\rho_{k}({\bm{x}})$ Distance to $k$-th nearest neighbor of $\bm{x}$} in $\mathcal{D}_{\mathcal{X}}$
$m_{\mathcal{L}}$ Regression model trained on $\mathcal{L}$
$l({\bm{x}},y,m_{\mathcal{L}})$ Error function measuring prediction error
Table 1: Table of notation used throughout the paper.
@@
Selecting points aiming to reduce the weighted fill distance of the training set is beneficial when traditional labeling methods, such as numerical simulations or laboratory experiments, are expensive or time-consuming. In such applications ML regression models provide fast label predictions for new data points. The quality of these predictions depends on the quality of the training data. Therefore, selecting an effective training set is crucial to ensure accurate predictions. Our research on minimizing the training set weighted fill distance can potentially identify sets that can enhance average prediction quality for various regression models and tasks, preventing the wastage of expensive labeling resources on training sets that may only benefit a specific learning model or task.
Acknowledgments
This work was partly supported by the German Federal Ministry of Education and Research (BMBF) within the projects MaGriDo (Mathematics for Machine Learning Methods for Graph-Based Data with Integrated Domain Knowledge), FKZ 05M20PDB and Rhine-Ruhr Center for Scientific Data Literacy (DKZ.2R), FKZ 16DKZ2030C.
@@ Appendix A The weighted fill distance is non-negative
The weighted fill distance defined in (2) is non-negative.
@@ Appendix B Assumptions
We recapture two assumptions from Climaco and Garcke (2024) that underlie our theoretical results. The first assumption concerns the data being analyzed and the relationship between data features and labels.
@@
With (17) we assume that the expected error on the training set is bounded. Moreover, with the Lipschitz continuity assumptions we limit our study to error functions that show a certain regularity. However, these regularity assumptions on the error function are not too restrictive and are connected with the regularity of the evaluated trained model as already shown in Climaco and Garcke (2024). For instance, the $\lambda_{l_{\mathcal{Y}}}$-Lipschitz regularity and the convexity in the second argument are verified by all $L_{p}$-norm error functions, with $1\leq p<\infty$. We also assume that $\mathbb{E}[|l({\bm{x}},Y,m_{\mathcal{L}})|\big|{\bm{x}}]<\infty$ for any given ${\bm{x}}\in\mathcal{X}$. This assumption formalizes the intuitive fact that, in applications, independently of the trained model and the feature vector considered, we can expect the error function to take finite values.
@@
Note that in (18) the expectation $\mathbb{E}\left[l({\bm{x}},Y,m_{\mathcal{L}})|{\bm{x}}\right]$ is independent of $p_{\mathcal{D}}$ and $p_{\mathcal{L}}$. This is because, as mentioned in Section 3, we consider a scenario where $p_{\mathcal{D}}$ may differ from $p_{\mathcal{L}}$ but the map connecting a data location ${\bm{x}}\in\mathcal{X}$ and its associated label value is independent on how the data is selected, that is, $p_{\mathcal{D}}\neq p_{\mathcal{L}}$ and $p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. In what follows we define $p(y|{\bm{x}}):=p_{\mathcal{D}}(y|{\bm{x}})=p_{\mathcal{L}}(y|{\bm{x}})$. Next, fixed ${\bm{X}}=\tilde{{\bm{x}}}\in\mathcal{X}$, we want to bound $\mathbb{E}\left[l(\tilde{{\bm{x}}},Y,m_{\mathcal{L}})|\tilde{{\bm{x}}}\right$]. To do that we use a result from Climaco and Garcke (2024): For fixed $\tilde{{\bm{x}}}\in\mathcal{X}$ and ${\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}$ we have
@@
$+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})\left(\int_{\mathcal{X}}p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}-\hskip-11.38092pt\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}{\int}\hskip-11.38092ptp_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})d{\bm{x}}\right)+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$\leq$ $\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}>0}{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}}{\int}\hskip-14.22636pt\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}C\left(1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}\right)p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})\left(1-\hskip-14.22636pt\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}{\int}\hskip-11.38092ptp_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})d{\bm{x}}\right)+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$\leq$ $\;C\thinspace\sup_{{\bm{x}}\in\mathcal{X}}\left(\min_{{\bm{x}}_{j}\in\mathcal{L}_{\mathcal{X}}}\|{\bm{x}}-{\bm{x}}_{j}\|_{2}\left(1-\frac{p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})}{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})}\right)\right)+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})\left(\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}{\int}\hskip-11.38092ptp_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})d{\bm{x}}\right)+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right]$
$=$ $\;CW_{\mathcal{L}_{\mathcal{X}}\negthinspace,\thinspace{\mathcal{X}}}\left(p_{\mathcal{X}_{\mathcal{L}}}||p_{\mathcal{X}_{\mathcal{D}}}\right)+(\lambda_{l_{\mathcal{Y}}}\epsilon+\epsilon_{\mathcal{L}})\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]+\mathbb{E}_{p_{\mathcal{L}}}\left[l({\bm{X}},Y,m_{\mathcal{L}})\right],$
where $\mathbb{P}_{\mathcal{L}}\left[p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}\right]:=\hskip-11.38092pt\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}<p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}{\int}\hskip-11.38092ptp_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})d{\bm{x}}=\int_{\mathcal{X}}p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\cdot\mathbf{1}\left\{p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})<p_{\mathcal{X}_{\mathcal{L}}}({\bm{x}})\right\}d{\bm{x}}$ is the probability, under $p_{\mathcal{X}_{\mathcal{L}}}$, that a randomly drawn sample is assigned higher likelihood by $p_{\mathcal{X}_{\mathcal{L}}}$ than by $p_{\mathcal{X}_{\mathcal{D}}}$. It quantifies the mismatch between $p_{\mathcal{X}_{\mathcal{L}}}$ and $p_{\mathcal{X}_{\mathcal{D}}}$. The last inequality follows from taking the supremum over ${\bm{x}}\in\mathcal{X}$, taking the constant terms out of the integral and noticing that $\hskip-11.38092pt\underset{\underset{p_{\mathcal{X}_{\mathcal{D}}}>0}{\underset{p_{\mathcal{X}_{\mathcal{D}}}\geq p_{\mathcal{X}_{\mathcal{L}}}}{\mathcal{X},}}}{\int}\hskip-11.38092ptp_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}\leq\underset{\mathcal{X}}{\int}p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})d{\bm{x}}=1$. ∎